# Program evaluation with remotely sensed outcomes

### v2

@@ Appendix E Discrete outcomes
    If Assumption 2.3(i) holds, then
    $\mathbb{E}\left[H_{d}(X,R)\frac{1\{D=d,S=e\}}{\Pr(D=d,S=e\mid X)}\mid X\right]=\mathbb{E}\left\{H_{d}(X,R)\Delta^{o}(d)\mid X\right\}\theta_{d}(X)$
    for $\Delta^{o}(d):=\left(\frac{1\{Y=y_{1},D=d,S=o\}}{\Pr(Y=y_{1},D=d,S=o)},\ldots,\frac{1\{Y=y_{|\mathcal{Y}|},D=d,S=o\}}{\Pr(Y=y_{|\mathcal{Y}|},D=d,S=o)}\right)$.
@@
    If Assumption 2.3(ii) holds, then
    $\mathbb{E}\left[H_{d}(X,R)\frac{1\{D=d,S=e\}}{\Pr(D=d,S=e\mid X)}\mid X\right]=\mathbb{E}\left\{H_{d}(X,R)\Delta^{o}\mid X\right\}\theta_{d}(X)$
    for $\Delta^{o}:=\left(\frac{1\{Y=y_{1},S=o\}}{\Pr(Y=y_{1},S=o)},\ldots,\frac{1\{Y=y_{|\mathcal{Y}|},S=o\}}{\Pr(Y=y_{|\mathcal{Y}|},S=o)}\right)$.
As in the main text, Theorem E.1 and Corollary E.1 imply that we can use existing results on conditional moment restrictions e.g. Chamberlain, (1987); Newey, (1993) for identification and inference on the average potential outcomes in the experimental sample. By further averaging over the conditional distribution $R\mid X$, we can construct unconditional moment restrictions using representations $H_{d}(X,R)$. Importantly, any choice of representation such that $\mathbb{E}\{H_{d}(X,R)\Delta^{o}(d)\mid X\}$ or $\mathbb{E}\{H_{d}(X,R)\Delta^{o}\mid X\}$ has full rank suffices for identification and valid asymptotic inference. The optimal representation is an obvious extension of our characterization in the main text.
As in the main text, Theorem E.1 and Corollary E.1 provide testable implications of the identifying assumptions through over-identifying restrictions: we can compare whether alternative choices of the representation $H_{d}$ yield significantly different estimates of the average potential outcomes in the experimental sample.
Appendix F Continuous outcomes
We extend our results outcomes with a continuous and bounded support.
Assumption F.1 (Bounded outcome).
Suppose that $Y(d)\in[-U,U]^{p}$ for some $U<\infty$. Suppose that for $d\in\{0,1\}$ and $s\in\{e,o\}$, $Y(d)\mid X,S=s,R$ has a positive density over $[-U,U]^{p}$ almost surely.
Let $f_{Y(d)\mid X,S,R}(y)$ denote the conditional density of $Y(d)\mid X,S=s,R$.
To extend our analysis to continuous outcomes, we construct bins over the support such that $y\in\mathcal{Y}_{\varepsilon}$ with $B_{y}(\varepsilon)$ defining an $l_{\infty}$-ball of radius $\varepsilon>0$ around the value $y$, meaning $||\widetilde{y}-\widetilde{y}^{\prime}||_{\infty}\leq\varepsilon$ for any two values $\widetilde{y},\widetilde{y}^{\prime}\in B_{y}(\varepsilon)$. Consequently, $\mathcal{Y}_{\varepsilon}$ defines an $\varepsilon$-cover of the support $[-U,U]^{p}$.
Let $\mathcal{Y}_{\varepsilon}=\{y_{1},\cdots,y_{|\mathcal{Y}_{\varepsilon}|}\}$ denote the support of the $\varepsilon$-cover. Furthermore, let
$\gamma_{y,d}(X,R)=\frac{\Pr\{Y\in B_{y}(\varepsilon),S=o,D=d\mid R,X\}}{\Pr\{Y\in B_{y}(\varepsilon),S=o,D=d\mid X\}}\mbox{ and }\gamma_{y}(X,R)=\frac{\Pr\{Y\in B_{y}(\varepsilon),S=o,\mid R,X\}}{\Pr\{Y\in B_{y}(\varepsilon),S=o,\mid X\}}.$
Proposition F.1 (Binning).
Suppose Assumptions 2.12.2, and F.1 hold.
  1. 1.
    If Assumption 2.3(i) holds, $\pi_{d}(X,R)=\sum_{y\in\mathcal{Y}_{e}}\gamma_{y,d}(X,R)\Pr\{Y(d)\in B_{y}(\epsilon)\mid S=e,X\}\mathrm{d}y.$
  2. 2.
    If Assumption 2.3(ii) holds, $\pi_{d}(X,R)=\sum_{y\in\mathcal{Y}_{e}}\gamma_{y}(X,R)\Pr\{Y(d)\in B_{y}(\epsilon)\mid S=e,X\}\mathrm{d}y.$
Proof.
The proof proceeds analogously to the proofs of Lemmas E.1 and  E.2, after replacing $Y=y$ events with $Y\in B_{y}(\varepsilon)$ events. Assumption F.1 and $\varepsilon>0$ guarantee that $Y\mid X,S,R\in B_{y}(\varepsilon)$ has a positive probability. ∎
Consequently, for any fixed $\epsilon>0$, we can estimate and conduct inference on the estimand
$\theta(\varepsilon):=\sum_{y\in\mathcal{Y}_{e}}y[\Pr\{Y(1)\in B_{y}(\epsilon)\mid S=e\}-\Pr\{Y(0)\in B_{y}(\epsilon)\mid S=e\}]$
$\theta:=\int_{[-U,U]^{p}}yf_{Y(1)\mid S=e}(y)\mathrm{d}y-\int_{[-U,U]^{p}}yf_{Y(0)\mid S=e}(y).$
Nonetheless, we expect that these two estimands will be close to each other for low-dimensional outcomes. When $p=1$, we show that the worst-case bias of $\theta(\varepsilon)$ is no greater than $\varepsilon>0$.
Proposition F.2 (Discretization bias).
Under the conditions of Proposition F.1, if $p=1$ then $|\theta-\theta(\varepsilon)\mid\leq\varepsilon$.
Proof.
To prove this result, let $f_{Y(d)\mid S}(y_{d})$ denote the conditional density of $Y(d)$ given $S$. As a first step, let us write
$\theta(\epsilon)$ $:=\sum_{\tilde{y}_{1}\in\mathcal{Y}_{e}}\tilde{y}_{1}\Pr\{Y(1)\in B_{\tilde{y}_{1}}(\epsilon)\mid S=e\}-\sum_{\tilde{y}_{0}\in\mathcal{Y}_{e}}\tilde{y}_{0}\Pr\{Y(0)\in B_{\tilde{y}_{0}}(\epsilon)\mid S=e\}$
$=\sum_{\tilde{y}_{1}\in\mathcal{Y}_{e}}\int_{B_{\tilde{y}_{1}}(\epsilon)}\tilde{y}_{1}f_{Y(1)\mid S=e}(y_{1})\mathrm{d}y_{1}-\sum_{\tilde{y}_{0}\in\mathcal{Y}_{e}}\int_{B_{\tilde{y}_{0}}(\epsilon)}\tilde{y}_{0}f_{Y(0)\mid S=e}(y_{0})\mathrm{d}y_{0},$
and analogously
$\theta$ $=\int y_{1}f_{Y(1)\mid S=e}(y_{1})\mathrm{d}y-\int y_{0}f_{Y(0)\mid S=e}(y_{0})$
$=\sum_{\tilde{y}_{1}\in\mathcal{Y}_{e}}\int_{B_{\tilde{y}_{1}}(\epsilon)}y_{1}f_{Y(1)\mid S=e}(y_{1})\mathrm{d}y_{1}-\sum_{\tilde{y}_{0}\in\mathcal{Y}_{e}}\int_{B_{\tilde{y}_{0}}(\epsilon)}y_{0}f_{Y(0)\mid S=e}(y_{0})\mathrm{d}y_{0}.$
Consequently, it follows that
$\left|\theta(\epsilon)-\theta\right|\overset{(1)}{\leq}\left|\sum_{\tilde{y}_{1}\in\mathcal{Y}_{e}}\int_{B_{\tilde{y}_{1}}(\epsilon)}\left(\tilde{y}_{1}-y_{1}\right)f_{Y(1)\mid S=e}(y_{1})\mathrm{d}y_{1}\right|+\left|\sum_{\tilde{y}_{0}\in\mathcal{Y}_{e}}\int_{B_{\tilde{y}_{0}}(\epsilon)}\left(\tilde{y}_{0}-y_{0}\right)f_{Y(0)\mid S=e}(y_{0})\mathrm{d}y_{0}\right|$
$\leq\max_{d\in\{0,1\}}2\left|\sum_{\tilde{y}_{d}\in\mathcal{Y}_{e}}\int_{B_{\tilde{y}_{d}}(\epsilon)}\left(\tilde{y}_{d}-y_{d}\right)f_{Y(d)\mid S=e}(y_{d})\mathrm{d}y_{d}\right|\overset{(2)}{\leq}\max_{d\in\{0,1\}}2\sum_{\tilde{y}_{d}\in\mathcal{Y}_{e}}\left|\int_{B_{\tilde{y}_{d}}(\epsilon)}\left(\tilde{y}_{d}-y_{d}\right)f_{Y(d)\mid S=e}(y_{d})\mathrm{d}y_{d}\right|$
[…]
where (1) and (2) use the triangle inequality, (3) follows by Hölder’s inequality, and (4) by the construction of the $\varepsilon$-cover. ∎
Appendix G Empirical application details
G.1 Real data from a randomized control trial
@@
Randomization took place at the level of the mandal, i.e. subdistrict, of Andhra Pradesh, India. Muralidharan et al., (2023) partition 396 mandals into the following subgroups:
@@
We study villages within mandals as the unit of analysis, where villages are defined by Asher et al., (2021). For each village, our treatment $D$ indicates whether the village received Smartcards in 2010. We interpret villages within the treated mandals as treated experimental units; villages within the buffer and untreated mandals as untreated experimental units; and villages within the non-study mandals as observational units with missing treatment. Finally, we drop villages with less than 100 individuals. This removes about 2% of villages.
@@
Table 5 summarizes village characteristics. We study $8,320$ villages, with an average population exceeding $2,000$ individuals per village. Villages are typically located in rural areas. The populated area within a typical village is geographically concentrated.
Table 5: Village summary statistics.
@@
Number of villages $2,260$ $853$ $2,931$ $2,276$
Average population $2,143$ $2,296$ $2,285$ $2,604$
Average fraction female $0.489$ $0.492$ $0.495$ $0.493$
Average fraction urban $0.002$ $0.001$ $0.003$ $0.004$
For each village, we collect poverty measurements to serve as the outcome. Following Asher and Novosad, (2020), the data sources are the 2012-2013 Social Economic and Caste Census (SECC), and the 2013 Indian Economic Census. Our main outcome variable, used in all three semi-synthetic exercises, indicates whether a village’s per capita consumption is in the bottom quartile. We consider two additional outcome variables in the second and third semi-synthetic exercises: does a village have only low income households, i.e. no earner making above 5,000 rupees; and does a village only low and middle income households, i.e. no earner making above 10,000 rupees. The definitions of low and middle income households is from the SECC.
G.2 Real satellite images
For each village, we extract satellite images to serve as the RSV. First, we extract coordinates for the perimeter of the village (Asher et al.,, 2021). Then, we extract luminosity from 2012 to 2020, summarized as a scalar in $\mathbb{R}$ (Asher et al.,, 2021). Finally, we extract satellite images from 2019, summarized as a high-dimensional, pre-trained embedding vector in $\mathbb{R}^{4000}$ (Rolf et al.,, 2021). The concatenation of these objects is our remotely sensed variable $R$.
In the first exercise, we truncate the satellite image vector to $\mathbb{R}^{1000}$ for computational tractability. In the second and third exercise, we use the full RSV.
@@ G.2 Real satellite images
Figure 11: No direct effects (Assumption 2.3(ii)) is plausible in the second and third exercises for the low consumption outcome. Within each plot, we compare $\Pr(R\mid S=s,D=0,Y=y)$ in the top row with $\Pr(R\mid S=s,D=1,Y=y)$ in the bottom row, using data from Muralidharan et al., (2023). Because the satellite image $R\in\mathbb{R}^{4000}$ is high dimensional, we visualize the density of its standardized first principal component.
While no direct effects (Assumption 2.3(ii)) holds by design in the first exercise, it must be defended in the second and third exercises. Figure 11 provides such evidence for the low consumption outcome. We compare $\Pr(R\mid S=s,D=0,Y=y)$ in the top row with $\Pr(R\mid S=s,D=1,Y=y)$ in the bottom row. The columns subset observations by $y\in\mathcal{Y}$, and the colors subset observations by $s\in\{e,o\}$. If, in a given column, the same-colored densities in the top and bottom row look similar, that is evidence that treatment only affects the RSV via the outcome. For example, focusing on the first column and the blue densities, we visually compare $\Pr(R\mid S=e,D=0,Y=0)$ in the top row with $\Pr(R\mid S=e,D=1,Y=0)$ in the bottom row. The densities are similar, as desired. Figure 13 provides similar evidence for the two other poverty outcomes.
@@
Figure 13: No direct effects (Assumption 2.3(ii)) is plausible in the second and third exercises for the additional poverty outcomes. Within each plot, we compare $\Pr(R\mid S=s,D=0,Y=y)$ in the top row with $\Pr(R\mid S=s,D=1,Y=y)$ in the bottom row, using data from Muralidharan et al., (2023). Because the satellite image $R\in\mathbb{R}^{4000}$ is high dimensional, we visualize the density of its standardized first principal component.
G.3 Implementation details
Across empirical exercises, we use random forest predictions: $\textsc{pred}_{Y}(R)$ estimates $\Pr(Y=1|S=o,R)$, $\textsc{pred}_{D}(R)$ estimates $\Pr(D=1|S=e,R)$, and $\textsc{pred}_{S}(R)$ estimates $\Pr(S=e|R)$. The first prediction appears in the common practice, while all three appear in the optimal representation.
In the first exercise, we use luminosity as well as the initial $1,000$ features of the satellite image embedding, for computational tractability of the $500$ replications. This truncation may be viewed as regularization bias, which we alleviate with cross fitting. This procedure is justified by Proposition 4.2.
In the second and third exercises, we use luminosity as well as all $4,000$ features of the satellite image embedding, since we are only running one replication. Random forests satisfy stability conditions, which allow us to eliminate cross fitting; the argument is a straightforward extension of Proposition 4.2, using stability in place of independence to handle the stochastic equicontinuity terms. See e.g. Chernozhukov et al., (2020, Theorem 2). Alternatively, we could use the limited complexity of random forests, along the lines of Chernozhukov et al., (2020, Theorem 3).
The technical details of the random forest implementation are as follows. We use the R package RandomForest with $100$ trees and the package default options. When predicting the outcomes, we set the class weights ten-to-one to balance the unbalanced labels of the outcome.
G.4 Comparing representations

### v3

@@ F.2 Deconvolution
    By equation (3),
    $f_{R}(r\mid S=e,X=x,D=d)=\int f_{R}(r\mid S=o,X=x,Y=y)f_{Y(d)}(y\mid S=e,X=x)\mathrm{d}y.$
@@
    By Bayes’ rule,
    $f_{R}(r\mid Y=y,S=o,X=x)$ $=\frac{f_{Y,S}(y,o\mid R=r,X=x)f_{R}(r|X=x)}{f_{Y,S}(y,o|X=x)}=\gamma_{y}^{0}(X,R)f_{R}(r|X=x),$
    $f_{R}\left(r\mid D=d,S=e,X=x\right)$ $=\frac{\Pr(D=d,S=e\mid R=r,X=x)f_{R}(r|X=x)}{\Pr(D=d,S=e|X=x)}=\pi_{d}(X,R)f_{R}(r|X=x).$
    Substituting these expressions into the previous step and canceling $f_{R}(r\mid X=x)$ yields the result.
Proposition F.3 identifies the conditional counterfactual density $f_{Y(d)}(y|S=e,X)$ as the solution to an integral equation. For example, consider the integral equation of Proposition F.3(ii):
$\pi_{d}(X,R)=\int_{y\in\mathcal{Y}}\gamma_{y}^{0}(X,R)f_{Y(d)}(y|S=e,X)\mathrm{d}y.$
The left hand side is a ratio of the regression functions $\Pr\left(D=d,S=e\mid X,R\right)$ and $\Pr\left(D=d,S=e\mid R\right)$, each of which is identified from the experimental sample. The right hand side is the composition of the integral operator $\tilde{f}\mapsto\int_{y\in\mathcal{Y}}\gamma_{y}^{0}(X,R)\tilde{f}(y|X)\mathrm{d}y$ and the target density $f_{Y(d)}(y|S=e,X)$. Clearly the integral operator is identified from the observational sample.
Recovering $f_{Y(d)}(y|S=e,X)$ is a classic statistical problem called a Fredholm integral equation of the first kind (Kress, 1989). In general, it is an ill posed inverse problem because to isolate $f_{Y(d)}(y|S=e,X)$ we must invert the integral operator. This problem is closely related to nonparametric instrumental variable regression and deconvolution, for which several nonparametric estimators are available in the literature; see e.g. Carrasco et al. (2007) for a review. Existence of the solution $f_{Y(d)}(y|S=e|X)$ will be guaranteed by Picard’s criterion (Kress, 1989). Uniqueness of the solution $f_{Y(d)}(y|S=e|X)$ will be guaranteed by a completeness condition on the integral operator (Newey and Powell, 2003). The rate of estimation for $f_{Y(d)}(y|S=e|X)$ will be the rate of estimation for $\pi_{d}(X,R)$, as throttled by the ill posedness of inverting the integral operator (e.g. Chen and Reiss, 2011). Articulating the technical details of such an approach is an exciting direction for future econometric theory.
Appendix G Crop Burning Illustration: Additional Details
We present details for the empirical example in Section 3.2, based on Jack et al. (2025).
The bias formula is derived in Appendix B. Within the main text, we take $\tilde{\beta}=1$ so that $\widetilde{\theta}=\beta\theta_{0}$ by Corollary B.1 since Jack et al. (2025) directly plug-in the RSV $R$ as the outcome in the experiment.
We classify a field as experimental ($S=e$) if it did not receive a random spot check. We classify a field as observational ($S=o$) if it received a random spot check. This example has complete cases (Assumption 3(ii)); the treatment varies in the observational sample.
In Section 3.2, we define $R\in\{0,1\}$ as the authors’ “maximum accuracy” classifier for whether a field has not been burned. Jack et al. (2025) construct another “balanced accuracy” classifier based on an alternative threshold rule. We find similar results in Table G.1, now defining $R\in\{0,1\}$ as the authors’ “balanced accuracy” classifier.
Table G.1: Underestimation of treatment effects in crop burning experiment: Revisited.
Common practice Bias Causal parameter
Estimand $\widetilde{\theta}$ $\beta$ $\theta$
Estimate 0.074 0.601 0.123
(0.049) (0.068) (0.086)
Notes: The RSV is the field-level “balanced accuracy” label defined by Jack et al. (2025), which applies a threshold rule to the predicted probability of not being burned. The “observational sample” has fields that received a random spot check, and the “experimental sample” has other fields. For illustration, we conduct linear estimation, controlling for stratum fixed effects. Standard errors are based on $5000$ bootstrap replications clustered at the village level.
Figure G.1 is provided purely for the sake of placing our empirical exercise of Section 3.2 on a map, keeping in mind that our analysis is at the field level.
(a) Experimental units only.
(b) Auxiliary sample: Observational units.
Figure G.1: We illustrate the samples in our re-analysis of the crop burning experiment in Jack et al. (2025), plotting the map of villages in Bathinda and Faridkot, which are two districts in Punjab, India. Experimental villages are those in which less than 50% of fields received a spot check. Observational villages are those in which more than 50% of fields received a spot check.
Appendix H Smartcard Illustration: Additional Details
H.1 Real Data from a Randomized Control Trial
@@
Randomization took place at the level of the mandal, i.e. subdistrict, of Andhra Pradesh, India. Muralidharan et al. (2023) partition 396 mandals into the following subgroups:
@@
We study villages within mandals as the units of analysis, where villages are defined by Asher et al. (2021). For each village, our treatment $D$ indicates whether the village received Smartcards in 2010. We interpret villages within the treated mandals as treated experimental units; villages within the buffer and untreated mandals as untreated experimental units; and villages within the non-study mandals as observational units with missing treatment. Finally, we drop villages with fewer than $100$ individuals. This removes about $2\%$ of villages.
@@
Table H.1 summarizes village characteristics. We study $8,320$ villages, with an average population exceeding $2,000$ individuals per village. Villages are located in rural areas. The populated area within a typical village is geographically concentrated.
Table H.1: Village summary statistics.
@@
Number of villages 2257 852 $29$ 2274
Average population 1985 2255 2105 2116
Average fraction female 0.491 0.492 0.496 0.497
For each village, we collect poverty measurements to serve as the outcome. Following Asher and Novosad (2020), the data sources are the 2012-2013 Socio-Economic and Caste Census (SECC), and the 2013 Indian Economic Census. Our main outcome variable, used in all three semi-synthetic exercises, indicates whether a village’s per capita consumption is in the bottom quartile. We consider two additional outcome variables in the second and third semi-synthetic exercises: does a village have only low income households, i.e. no earner making above 5,000 rupees; and does a village only low and middle income households, i.e. no earner making above 10,000 rupees. The definitions of low and middle income households is from the SECC.
H.2 Real Satellite Images
For each village, we extract satellite images to serve as the RSV. First, we extract coordinates for the perimeter of the village (Asher et al., 2021). Then, we extract luminosity measures, a vector in $\mathbb{R}^{50}$, which includes the minimum, maximum, mean, and sum of night light within each polygon, along with the total number of pixels in the polygon, from 2012 to 2020 (Asher et al., 2021). Finally, we extract satellite images from 2019, summarized as a high-dimensional, pre-trained embedding vector in $\mathbb{R}^{4000}$ (Rolf et al., 2021). The concatenation of these objects is our remotely sensed variable $R$.
In the first exercise, we truncate the RSV from $\mathbb{R}^{4050}$ to $\mathbb{R}^{1050}$ for computational tractability. In particular, we use the initial $1,000$ features of the pre-trained embedding vector. In the second and third exercise, we use the full RSV in $\mathbb{R}^{4050}$ .
@@ H.2 Real Satellite Images
Figure H.1: No direct effects (Assumption 3(ii)) is plausible in the second and third exercises for the low consumption outcome. Within each plot, we compare $f_{R}(R\mid S=s,D=0,Y=y)$ in the top row with $f_{R}(R\mid S=s,D=1,Y=y)$ in the bottom row, using data from Muralidharan et al. (2023). Because the RSV $R$ is high-dimensional, we visualize the density of its standardized first principal component.
While no direct effects (Assumption 3(ii)) holds by design in the first exercise, it must be defended in the second and third exercises. Figure H.1 provides such evidence for the low consumption outcome. We compare $f_{R}(R\mid S=s,D=0,Y=y)$ in the top row with $f_{R}(R\mid S=s,D=1,Y=y)$ in the bottom row. The columns subset observations by $y\in\mathcal{Y}$, and the colors subset observations by $s\in\{e,o\}$. If, in a given column, the same-colored densities in the top and bottom row look similar, that is evidence that treatment only affects the RSV via the outcome. For example, focusing on the first column and the blue densities, we visually compare $f_{R}(R\mid S=e,D=0,Y=0)$ in the top row with $f_{R}(R\mid S=e,D=1,Y=0)$ in the bottom row. The densities are similar, as desired. Figure H.3 provides similar evidence for the two other poverty outcomes.
@@
Figure H.3: No direct effects (Assumption 3(ii)) is plausible in the second and third exercises for the additional poverty outcomes. Within each plot, we compare $f_{R}(R\mid S=s,D=0,Y=y)$ in the top row with $f_{R}(R\mid S=s,D=1,Y=y)$ in the bottom row, using data from Muralidharan et al. (2023). Because the RSV $R$ is high-dimensional, we visualize the density of its standardized first principal component.
H.3 Implementation Details
Across empirical exercises, we use random forest predictions: $\textsc{pred}_{Y}(R)$ estimates $\Pr(Y=1|S=o,R)$, $\textsc{pred}_{D}(R)$ estimates $\Pr(D=1|S=e,R)$, and $\textsc{pred}_{S}(R)$ estimates $\Pr(S=e|R)$. The first prediction appears in the common practice, while all three appear in the efficient representation.
In the first exercise, we use luminosity as well as the initial $1,000$ features of the satellite image embedding, for computational tractability of the $500$ replications. Using this subset of features makes cross-fitting computationally feasible. This procedure is justified by Proposition 3.
In the second and third exercises, we use luminosity measures as well as all $4,000$ features of the satellite image embedding, since we are only running one replication. Random forests satisfy stability conditions, which allow us to eliminate cross-fitting; the argument is a straightforward extension of Proposition 3, using stability in place of independence to handle the stochastic equicontinuity terms. See e.g. Chernozhukov et al. (2020, Theorem 2). Alternatively, we could use the limited complexity of random forests, along the lines of Chernozhukov et al. (2020, Theorem 3).
The technical details of the random forest implementation are as follows. We use the R package randomForest with $100$ trees and the package default options. When predicting the outcomes, we set the class weights ten-to-one because one outcome value is much more frequent than the other.
H.4 Comparing Alternative Representations